OpenStudy (anonymous):

Someone please help me!! I'll give out medals!! Show the steps for the integration of: (x+3)/((x^3)+1)^(1/2)

OpenStudy (anonymous):

\[\int\limits_{?}^{?} \frac{ x+3 }{ \sqrt{x^3+1} }\]

OpenStudy (anonymous):

How do I integrate this?

geerky42 (geerky42):

Exactly what are upper and lower limits?

OpenStudy (daniel14):

hmm

geerky42 (geerky42):

Or is it supposed to be indefinite integral?

OpenStudy (anonymous):

indefinite integral...

geerky42 (geerky42):

Well, it won't look nice... http://www.wolframalpha.com/input/?i=integral+%28x%2B3%29%2F%28sqrt%28x%5E3%2B1%29%29